OpenStudy (jpsmarinho):

If f(x) = x^3+2x+1 and g is the inverse function of f, so, g'(1) is equal to: -1/4 -1/2 0 1/4 1/2 Someone know how to do?

myininaya (myininaya):

\[g=f^{-1}\] \[(f^{-1})'(1)=\frac{1}{f'(f^{-1}(1))}\]

myininaya (myininaya):

Do you know how to find \[f^{-1}(1)=? \]

OpenStudy (jpsmarinho):

I've found that f inverse is equals to x-1/(y^2+2). Is this correct?

myininaya (myininaya):

no you can't easily find the general form of the inverse answer my question :)

myininaya (myininaya):

\[f^{-1}(0)=? \text{ implies } f(?)=0\] What can you plug into your function that will give you 0 as output?

myininaya (myininaya):

oops that 0 is suppose to be a 1 by the way

myininaya (myininaya):

We are looking for \[f^{-1}(1)=? \text{ \implies } f(?)=1 \]

OpenStudy (jpsmarinho):

Sorry, I went dinner. ^^. It makes sense! Then x^2 = -1/3 right?

myininaya (myininaya):

answer my question first I don't know what you mean what can you plug into your function so that the output is 1

myininaya (myininaya):

what does \[f(0)=?\]

OpenStudy (jpsmarinho):

I don't know ^^. This is the enuncioation of the exercis!

myininaya (myininaya):

\[f(x)=x^3+2x+1\] To evaluate f(0) just replace all the x's with 0

myininaya (myininaya):

Tell what is \[f(0)=0^3+2(0)+1\]

OpenStudy (jpsmarinho):

But the derivative of f(x) that equal to 1 right?

myininaya (myininaya):

? Right now we are trying to find the ? in f^{-1}(1)=? \text{ \implies } f(?)=1

myininaya (myininaya):

\[f^{-1}(1)=? \text{ \implies } f(?)=1\]

OpenStudy (jpsmarinho):

Aha right

myininaya (myininaya):

Can you evalue f(0) for me ?

myininaya (myininaya):

Like I wrote what f(0) was all you have to do is simplify

OpenStudy (jpsmarinho):

f(0) = 1

myininaya (myininaya):

great so remember we are trying to evaluate \[(f^{-1})'(1)=\frac{1}{f'(f^{-1}(1))}=\frac{1}{f'(0)}\] Now you need find f' and evaluate f'(0)

OpenStudy (jpsmarinho):

Is equal to 2 right?

myininaya (myininaya):

first what did you get for f'

OpenStudy (jpsmarinho):

3x^2 +2

myininaya (myininaya):

Ok great

OpenStudy (jpsmarinho):

The answer is 1/2 right?

myininaya (myininaya):

yes

OpenStudy (jpsmarinho):

Thanks @myininaya !!